OpenStudy (anonymous):

d=vt+1/2at^2 solve for a

OpenStudy (john_es):

\[a=2(d-vt)/t^2\]

OpenStudy (anonymous):

\[d = vt + \frac{ 1 }{ 2 }a t^{2}\] Step 1) Subtract vt from both sides \[d - vt = \frac{ 1 }{ 2 }a t^{2}\] Step 2) Multiply both sides by 2 \[2(d - vt) = a t^{2}\] Step 3) Divide both sides by t-squared \[\frac{2(d - vt)}{ t^{2}} = a\]